OpenStudy (anonymous):

Precalculus help please? I'm really confused on what to do. How do you solve for x? sec (3x/2) = -2

OpenStudy (anonymous):

start with \[\cos(\frac{3x}{2})=-\frac{1}{2}\] solve that for \(\frac{3x}{2}\) and then solve for \(x\)

OpenStudy (anonymous):

cos (3x/2) = -1/2 = cos 2Pi/3 and cos 4Pi/4. 1). 3x/2 = 2Pi/3 -> x =..... 2). 3x/2 = 4Pi/3 -> x = ....

OpenStudy (anonymous):

Oh, I'm doing the arccos(-1/2) wrong...I kind of forgot how to do it. Can you show me the steps please?

OpenStudy (anonymous):

sec(3x/2) = -2. Since cos x =1/sec x, then cos (3x/2) = -1/2 . Trig Conversion Table, or calculator give -1/2 = cos (2Pi/3) and = cos (4Pi/3). Theses 2 arcs have the same cos value (-1/2) There are 2 solutions: 1. 3x/2 = 2Pi/3 -> x = ... 2. 3x/2 = 4Pi/3 -> x = ...